OpenStudy (anonymous):

evaluate -4(a+b)-10a/b when a=8 and b=5

OpenStudy (anonymous):

Simplify the following: -4 (8+5)-10×8/5 Evaluate 8+5. 8+5 = 13: -413-10×8/5 Multiply -4 and 13 together. -4×13 = -52: -52-10×8/5 Multiply -10 and 8 together. -10×8 = -80: -80/5-52 Reduce (-80)/5 to lowest terms. The gcd of -80 and 5 is 5, so (-80)/5 = (5 (-16))/(5×1) = 5/5×-16 = -16: -16-52 Group the negative terms in -52-16 together and factor out the minus sign. -52-16 = -(52+16): -(52+16) Evaluate 52+16 using long addition. | 5 | 2 + | 1 | 6 | 6 | 8: Answer: | | -68

OpenStudy (anonymous):

-4(a+b) - (10a)/b = -4(8+5) - (10x8)/5 = (-4x13) - (80/5) = -52 -16 = -68

OpenStudy (anonymous):

First get an easyer problem from this way :! first : -4(a+b) = -4a-4b Then :-10a/b = -10a/b Now we have : -4a-4b-10a/b Now give the numbers : -4(8)-4(5)-10(8)/5 So we have : -32-20-80/5 : -32-20-16 :-4 Understond ?